OpenStudy (roadjester):

How do I solve this isobaric differential equation? \[\large ydx + (2x^2y+x)dy=0\]

OpenStudy (roadjester):

\[\large ydx + (2x^2y+x)dy=0\]

OpenStudy (sirm3d):

this is bernoulli equation in the dependent variable x. \[yx'+x=-2x^2y\\x'+\frac{1}{y}x=-2x^2\]